# A highly accurate procedure for computing globally optimal Wannier functions in one-dimensional crystalline insulators

### v3

@@
A standard task in solid state physics and quantum chemistry is the computation of localized molecular orbitals known as Wannier functions. In this manuscript, we propose a new procedure for computing Wannier functions in one-dimensional crystalline materials. Our approach proceeds by first performing parallel transport of the Bloch functions using numerical integration. Then a simple analytically computable correction is introduced to yield the optimally localized Wannier function. The resulting scheme is rapidly convergent and proven to produce globally optimal Wannier functions. The analysis in this manuscript can also be viewed as a proof of the existence of exponentially localized Wannier functions in one dimension. We illustrate the performance of the scheme by a number of numerical experiments.
@@
Abinand $\mbox{}^{\dagger}$, Hanwen Zhang $\mbox{}^{\ddagger}$
September 23, 2024
$\mbox{}^{\dagger}$ Department of Mathematics, UC Davis, Davis, CA 95616
$\mbox{}^{\ddagger}$ Department of Mathematics, Yale University, New Haven, CT 06511
@@ 1 Introduction
Wannier functions are localized molecular orbitals in crystalline materials. They provide a convenient basis for many tasks in electronic structure calculations, such as when using the tight-binding approximation and understanding the conductivity and polarization properties of certain materials (cf. [20]). As such, the computation of Wannier functions is of interest in the development of new materials, such as semiconductors, solar cells, and topological insulators.
@@
From the point of view of mathematics, the problem amounts to solving a series of parameter-dependent eigenvalue problems, such that the eigenvectors are smooth as a function of the parameter. The most common approach is to minimize the Marzari–Vanderbilt functional using some form of gradient descent with high quality initial guesses computed by density-matrix based methods [6, 7, 16]. In this manuscript, we propose an alternative approach that requires no iterative optimization. Instead, we solve the parallel transport problem via numerical integration to determine the Bloch functions and band functions. We then show that a phase correction yields the maximally-localized Wannier function. The analysis in this paper for exponential localization is based on Kato’s analytic perturbation theory and is related to the work [5]. The identification of the global optimal solution is a simple consequence of the analysis.
@@
The algorithm in this paper is closely related to the one dimensional case in [4, 15]. However, this approach differs from these existing works in three crucial aspects. First, we provide a complete numerical procedure that starts with the Schrödinger equation and ends with the evaluation of the Wannier function. Second, this approach is high-order convergent and amendable to further improvement by switching to arbitrarily higher-order ODE solvers. Finally and most significantly, we are able to directly compute a provably maximally localized Wannier function without any type of iterative optimization. We refer the reader to Remark 4.1 for further details.
@@
This is the first in a series of papers; in subsequent papers we will consider the multiband case in one dimension, and single and multiband cases in higher dimensions.
@@ 2.1 Notation
We let $\ell^{2}(\mathbb{Z})$ denote the Hilbert space for complex sequences $\mathbf{x}=\left\{x_{i}\right\}_{i=-\infty}^{\infty}$ such that $\sum_{i=-\infty}^{\infty}\absolutevalue{x_{i}}^{2}<\infty$ with the standard inner product. We denote the right shift operator in $\ell^{2}(\mathbb{Z})$ by $R$, where
@@
We choose the principal branch of the $\log$ function so that we assume all real phases $\phi$ in $e^{\mathrm{i}\phi}$ take value in $(-\pi,\pi]$.
@@ 2.2 Bloch functions
We refer to the interval $[-a/2,a/2)$ as the primitive unit cell:
@@
as the periodic for the reciprocal lattice
@@
The interval $[-\Omega/2,\Omega/2)$ is usually called the first Brillouin zone (BZ), and we denote the interval by $I_{\mathrm{bz}}:$
@@
The following lemma, stating that the operator defining (13) is analytic in $k$ and has a discrete spectrum for piecewise continuous potential $V$ in $I_{\mathrm{uc}}$, can be found in [17].
@@ 2.3 Wannier functions
Wannier functions are highly non-unique since the Bloch functions in (9) with the normalization condition in (16) is only determined up to a complex number of unit modulus. More explicitly, for any $\psi_{k}^{(j)}$ satisfying (9), so will any $e^{-\mathrm{i}\phi(k)}\psi_{k}^{(j)}$ with a real $\phi(k)$ for $k\in I_{\mathrm{bz}}$ without affecting the normalization (16). In this paper, we refer to numbers of the form $e^{-\mathrm{i}\phi}$ as phase factors. For both physical and computational purposes, it is highly desirable to choose $\psi_{k}^{(j)}$ so that the Wannier functions are well-localized.
@@ 2.4 Analyticity of eigenvalues and eigenvectors
The above theorem is still applicable when the eigenvalues become degenerate. In this case, the eigenprojectors at degenerate points must be viewed as limits of those approaching the degenerate points. We refer the reader to Theorem 1.10 and Remark 1.11 in [13] for details.
Remark 2.5.
Theorem 2.3 concerns only the eigenprojectors, not the eigenvectors; this is because for any an eigenvector $\mathbf{v}_{i}(z)$ of $T(z)$, one can multiply the vector by a $z$-dependent phase factor $e^{-\mathrm{i}\phi(z)}$ ($\phi(z)$ is real for real $z$) while the new vector $e^{-\mathrm{i}\phi(z)}\mathbf{v}_{i}(z)$ is still an eigenvector. The phase factor could change the smoothness of $\mathbf{v}_{i}(z)$ while the projector $P_{i}(z)=\mathbf{v}_{i}(z)\mathbf{v}^{*}_{i}(z)$ is unchanged for real $z$.
@@
While the above results only apply to self-adjoint analytic family of finite-dimensional matrices $T(z)$ in $D$, they straightforwardly generalize to a class of infinite-dimensional operators in Hilbert space. The definition for the analytic family $T(z)$ being self-adjoint in a domain $D$ intersecting the real axes is unchanged. To extend Theorems 2.3 and 2.6 to the operator case Kato defined a class of operators where a subset of their spectra can be separated by a simple close curve in the complex plane from the rest.
Kato showed that, for a self-adjoint analytic family $T(z)$ possessing a finite set of isolated spectrum that can be separated for $z\in D$, the finite-dimensional version Theorems 2.3 and 2.6 can be applied to the finite-dimensional subspace corresponding to the separated eigenvalues. We summarize this extension in the following theorem and refer the reader to Sections 7.1.3 and 7.3.1 in [13] for details.
@@
Suppose that $T(z)$ is an analytic family of operators in Hilbert space for $z\in D$. Suppose further that $T(z)$ is self-adjoint, defined analogously as the matrix case in Theorem 2.3. Theorems 2.3 and 2.6 also apply to any finite systems of eigenvalues, eigenprojectors and eigenvectors of $T(z)$, where the eigenvalues can be separated from the rest of the spectrum of $T(z)$.
@@ 2.5 Derivatives of eigenvalues and eigenvectors
Remark 2.9.
In the physics literature, the formula (28) is usually written using the spectral decomposition of $T$ in the following form
@@ 3.1 Analysis of Bloch functions
We observe that Theorem 2.7 does not guarantee the function $E(k)$ or $u(x,k)$ to be periodically analytic in $k\in I_{\mathrm{bz}}$, which will be shown to be necessary for exponentially localized Wannier functions. If $E(k)$ were to become degenerate for some $k_{0}\in I_{\mathrm{bz}}$, it could go to another branch so that it does not return to $E(-\Omega/2)$ when it reaches $E(\Omega/2)$. When $E$ is not degenerate, $E(k)$ will be periodically analytic in $k\in I_{\mathrm{bz}}$. For the eigenvector $u(x,k)$, as discussed in Remark 2.5, whenever $u(x,k)$ is analytic in $k\in I_{\mathrm{bz}}$, so is any $e^{-\mathrm{i}\phi(k)}u(x,k)$ with a real analytic $\phi$, which need not be periodic in $I_{\mathrm{bz}}$.
@@
$\sum_{m=-\infty}^{\infty}\absolutevalue{a_{m}(k)}^{2}=\frac{1}{a}.$ (37)
For simplicity, we assume for the rest of the paper that $u(x,k)$ as a function of $x$ is a periodic function in $C^{1}(I_{\mathrm{uc}};\mathbb{C})$ and $\partial_{x}u(x,k)$ has bounded variation for all $k\in I_{\mathrm{bz}}$, so that the Fourier coefficients $a_{m}(k)$ decays as |$\absolutevalue{m}^{-2}$ for large $m$ [23, Section 2.6]. Thus, we have a positive constant $D$ independent of $k$, such that
$|a_{m}(k)|\leq\frac{D}{|m|^{2}},\quad m\in\mathbb{Z}.$ (38)
@@
Suppose $u(x,k)$ is chosen to be analytic in $k\in I_{\mathrm{bz}}$. Its Fourier coefficients $a_{m}(k)$ in (35) are analytic for $k\in I_{\mathrm{bz}}$.
@@
Lemma 39 shows that we can obtain the function $\alpha$ defined in $\mathbb{R}$ by “unfolding” the Fourier coefficients $a_{m}(k)$ defined in $I_{\mathrm{bz}}$ for $m\in\mathbb{Z}$ such that $a_{m}(k)$ takes up the interval $[-\Omega/2+m\Omega,\Omega/2+m\Omega)$. We summarize this fact in the following lemma, together with Lemma 3.3 and the decay condition (38) in terms of $\alpha$.
@@
where $\alpha$ is defined in (39). The function $\alpha(k)$ can be chosen to be analytic in the intervals $[-\Omega/2+m\Omega,\Omega/2+m\Omega)$ for $m\in\mathbb{Z}$. Furthermore, |$\absolutevalue{\alpha(k)}$ | decays no slower than $|k|^{-2}$ as $|k|\to\infty$.
@@
Since $u(x,k)$ may not be periodically analytic in $k\in I_{\mathrm{bz}}$ (see Remark 3.2), the function $\alpha$ can be a discontinuous function in $\mathbb{R}$. Lemma 3.3 only ensures $\alpha$ is analytic in each interval $[-\Omega/2+m\Omega,\Omega/2+m\Omega)$ for $m\in\mathbb{Z}$. Moreover, the function $\alpha$ is formed from a very special choice of coefficients $a_{m}$ since the corresponding $u(x,k)$ is chosen to be analytic in $k$; this does not happen if one computes $u(x,k)$ at each point $k$ independently (see Remark 2.5) so that $u(x,k)$ is an arbitrarily irregular function in $k$.
@@ 3.2 Properties of Wannier functions
We observe that (47) only hold formally since $\alpha$ may not even be integrable (see Remark 3.6). However, whenever $\alpha$ is chosen to be integrable, (47) shows that $\frac{2\pi}{\Omega}\alpha$ is the Fourier transform of $W_{0}$ in (46) and we have the following theorem.
@@
Suppose $u$ in Lemma 3.5 is chosen so that the function $\alpha$ is integrable. The corresponding Wannier function $W_{0}$ in (46) is well defined and is given by the formula
@@
Its Fourier transform denoted by $\widehat{W}_{0}$ is given by the formula
@@
By the Riemann–Lebesgue lemma, |$\absolutevalue{W_{0}(x)}$ | goes to zero as |$\absolutevalue{x}\to\infty$ | → ∞.
Thus, if $\alpha$ is chosen to have poor regularity, the Wannier function $W_{0}$ is poorly localized. To define its variance $\alpha$ needs to be smooth so that |$\absolutevalue{W_{0}(x)}$ | goes to zero faster than $1/\absolutevalue{x}$ | as |$\absolutevalue{x}\to\infty$ | → ∞. Next, we assume that $\alpha$ is chosen so that its derivative is integrable. The following lemma contains the formulas for the first and second moments of $W_{0}$ in terms of $\alpha$ for defining its variance. They are consequences of the one-dimensional case of (7) and (8) in [15], first derived in [2].
@@
Suppose that $u$ in Lemma 3.5 is chosen so that the function $\alpha$ has an integrable derivative $\alpha^{\prime}$ in $\mathbb{R}$. We have the following formulas for the first and second moments of $W_{0}$:

### v4

@@
A standard task in solid state physics and quantum chemistry is the computation of localized molecular orbitals known as Wannier functions. In this manuscript, we propose a new procedure for computing Wannier functions in one-dimensional crystalline materials. Our approach proceeds by first performing parallel transport of the Bloch functions using numerical integration. Then a simple analytically computable correction is introduced to yield the optimally localized Wannier function. The resulting scheme is rapidly convergent and is proven to yield real-valued Wannier functions that achieve global optimality. The analysis in this manuscript can also be viewed as a proof of the existence of exponentially localized Wannier functions in one dimension. We illustrate the performance of the scheme by a number of numerical experiments.
@@
[…]
October 17, 2025
[…]
@@ 1 Introduction
The eigenfunctions of the Schrödinger equation form a natural basis for electronic structure calculations. These eigenfunctions provide a complete description of the quantum states of a system and are the foundation for analyzing spectral properties, energies, and response to external perturbations. For atoms and molecules in isolation, the eigenfunctions are spatially localized, reflecting the confinement of electrons to specific regions of space, and they lend themselves to intuitive chemical and physical interpretations. In crystalline materials, on the other hand, the situation changes fundamentally. The periodicity of the potential leads, by Bloch’s theorem, to eigenfunctions (known in this case as Bloch functions) that are not localized in real space but instead exhibit quasiperiodicity. The delocalized nature of the eigenfunctions makes them less useful in electronic structure calculations. More useful in this setting are Wannier functions which are localized molecular orbitals in crystalline materials. These then can be used for the tight-binding approximation and to understand the conductivity and polarization properties of certain materials (cf. [22]). As such, the computation of Wannier functions is of interest in the development of new materials, such as semiconductors, solar cells, and topological insulators.
Formally, Wannier functions are obtained as the Fourier transform of Bloch functions with respect to crystal momentum. However, the Bloch functions are defined only up to a momentum-dependent phase (or, more generally, unitary transformation), often referred to in physics as a “gauge.” Different choices of gauge yield different Wannier functions, with the most relevant for applications being the optimally localized Wannier functions. Mathematically, the construction amounts to solving a family of parameter-dependent eigenvalue problems while ensuring that the eigenvectors vary smoothly with respect to the momentum parameter.
The most common approach is to first compute the Bloch functions and then choose the gauge to minimize the variance of the resulting Wannier function (known as the Marzari–Vanderbilt functional) using some form of gradient descent with high quality initial guesses computed by density-matrix based methods [6, 7, 18]. Unique to the one-dimensional case, it is also known that the globally optimally localized Wannier functions are the eigenfunctions of the projected position operator [17, 22].
@@
In this manuscript, we propose an alternative approach that requires no iterative optimization and is generalizable to higher dimensions. We solve a parallel transport problem via numerical integration to determine the Bloch functions and band functions. We then show that a phase correction yields real-valued Wannier functions that achieve global optimality, thereby eliminating the need for iterative optimization or for solving the eigenvalue problem of the projected position operator. We also provide complete analysis for the construction, establishing for the first time a constructive proof for the existence of exponentially localized Wannier function in one dimension possessing global optimality via parallel transport. The analysis in this paper for exponential localization is based on Kato’s analytic perturbation theory and is related to the work [5]. The identification of the global optimal solution is a simple consequence of the analysis.
@@
The algorithm in this paper is closely related to the one dimensional case in [4, 16, 22], which are also based on parallel transport. However, our approach differs from these existing works in three crucial aspects. First, we provide a complete numerical procedure that starts with the Schrödinger equation and ends with the evaluation of the Wannier function, where we solve the parallel transport equation exactly up to discretization errors only. Second, this convergence of our approach is of arbitrarily high-order with the order determined solely by choice of integrator. This is in contrast to the approach of [4, 16, 22] which has a parallel transport procedure that is fundamentally limited to low order. Finally and most significantly, we are able to directly compute a provably globally optimal Wannier function without any type of iterative optimization.
Furthermore, we point out that our analysis shows the parallel transport scheme in one dimension in [16, 22] (see Remark 4.3), although only a second-order approximation, also yields exponentially localized Wannier functions when combined with the same correction procedure in this paper. To the best of our knowledge, this has long been a heuristic among physicists [15, 22], but has never been rigorously proven. The resulting Wannier functions from this scheme, however, differ slightly from the globally optimal ones presented in this work, with the discrepancy arising from the approximation in the parallel transport step. This observation is particularly relevant in settings where it is advantageous to first obtain the Bloch functions, such as in density functional theory. To avoid digression, we will report a detailed analysis of the above points at a later date.
@@
This is the first in a series of papers; the extension of our results to single band matrix models in two dimensions can be found in the preprint [24]. In subsequent papers we will consider the multiband case both in one dimension and higher dimensions. We will briefly mention the challenges for these extensions in Section 7.
@@ 2.1 Notation
[…]
@@
We choose the principal branch of the $\log$ function so that we assume all real phases $\varphi$ in $e^{\mathrm{i}\varphi}$ take value in $(-\pi,\pi]$.
@@ 2.2 Bloch functions
We refer to the interval [a/2,a/2)[-a/2,a/2) as the primitive unit cell:
@@
as the period for the reciprocal lattice
@@
The interval $\Omega/2,\Omega/2)$ is usually called the first Brillouin zone (BZ), and we denote the interval by $I_{\mathrm{bz}}:$
@@
The following result in [19] asserts that, if $V$ is piecewise continuous, then the operator associated with (13) is analytic in $k$ in the standard sense (cf. [13, 19]); that is, its resolvent depends analytically on $k$ as a bounded operator. Furthermore, the operator has purely discrete spectrum.
@@ 2.3 Wannier functions
Wannier functions are highly non-unique since the Bloch functions in (9) with the normalization condition in (16) are only determined up to a complex number of unit modulus. More explicitly, if any $\psi_{k}^{(j)}$ satisfies (9), so will any $e^{-\mathrm{i}\varphi(k)}\psi_{k}^{(j)}$ with a real $\varphi(k)$ for $k\in I_{\mathrm{bz}}$ without affecting the normalization (16). In this paper, we refer to numbers of the form $e^{-\mathrm{i}\varphi}$ as phase factors. For both physical and computational purposes, it is highly desirable to choose $\psi_{k}^{(j)}$ so that the Wannier functions are well-localized.
@@ 2.4 Analyticity of eigenvalues and eigenvectors
It is worth noting that the above theorem is still applicable when the eigenvalues become degenerate. In this case, the eigenprojectors at degenerate points must be viewed as limits of those approaching the degenerate points. We refer the reader to Theorem 1.10 and Remark 1.11 in [13] for details.
@@
Theorem 2.3 concerns only the eigenprojectors, not the eigenvectors; this is because for any an eigenvector $\mathbf{v}_{i}(z)$ of $T(z)$, one can multiply the vector by a $z$-dependent phase factor $e^{-\mathrm{i}\varphi(z)}$ ($\varphi(z)$ is real for real $z$) while the new vector $e^{-\mathrm{i}\varphi(z)}\mathbf{v}_{i}(z)$ is still an eigenvector. The phase factor could change the smoothness of $\mathbf{v}_{i}(z)$ while the projector $P_{i}(z)=\mathbf{v}_{i}(z)\mathbf{v}^{*}_{i}(z)$ is unchanged for real $z$.
@@
While the above results only apply to self-adjoint analytic family of finite-dimensional matrices $T(z)$ in $D$, they straightforwardly generalize to a class of infinite-dimensional operators in Hilbert space. The definition for the analytic family $T(z)$ being self-adjoint in a domain $D$ intersecting the real axes is unchanged. To extend Theorems 2.3 and 2.5 to the operator case Kato defined a class of operators where a subset of their spectra can be separated by a simple close curve in the complex plane from the rest.
Kato showed that, for a self-adjoint analytic family $T(z)$ possessing a finite set of isolated spectrum that can be separated for $z\in D$, the finite-dimensional version Theorems 2.3 and 2.5 can be applied to the finite-dimensional subspace corresponding to the separated eigenvalues. We summarize this extension in the following theorem and refer the reader to Sections 7.1.3 and 7.3.1 in [13] for details.
@@
Suppose that $T(z)$ is an analytic family of operators in Hilbert space for $z\in D$. Suppose further that $T(z)$ is self-adjoint, defined analogously as the matrix case in Theorem 2.3. Theorems 2.3 and 2.5 also apply to any finite systems of eigenvalues, eigenprojectors and eigenvectors of $T(z)$, where the eigenvalues can be separated from the rest of the spectrum of $T(z)$.
@@ 2.5 Derivatives of eigenvalues and eigenvectors
We note that in the physics literature, the formula (28) is usually written using the spectral decomposition of $T$ in the following form
@@ 3.1 Analysis of Bloch functions
We observe that Theorem 2.6 does not guarantee the function $E(k)$ to be periodically analytic in $k\in I_{\mathrm{bz}}$ or $u(x,k)$ to satisfy (34), which will be shown to be necessary for exponentially localized Wannier functions. If $E(k)$ were to become degenerate for some $k_{0}\in I_{\mathrm{bz}}$, it could go to another branch so that it does not return to $E(-\Omega/2)$ when it reaches $E(\Omega/2)$. When $E$ is not degenerate, $E(k)$ will be periodically analytic in $k\in I_{\mathrm{bz}}$. For the eigenvector $u(x,k)$, as discussed in Remark 2.4, whenever $u(x,k)$ is analytic in $k\in I_{\mathrm{bz}}$, so is any $e^{-\mathrm{i}\varphi(k)}u(x,k)$ with a real analytic $\varphi$, which need to satisfy (34).
@@
[…]
$|.|a_{m}(k)|\leq\frac{D}{|m|^{2}},\quad\mbox{as $\absolutevalue{m}\rightarrow\infty$}\,$. (38)
@@
Suppose $u(x,k)$ is chosen to be analytic in $k\in I_{\mathrm{bz}}$. Its Fourier coefficients $a_{m}(k)$ in (35) are analytic in $k\in I_{\mathrm{bz}}$.
@@
Lemma 39 shows that we can obtain the function $\alpha$ defined on $\mathbb{R}$ by “unfolding” the Fourier coefficients $a_{m}(k)$ defined on $I_{\mathrm{bz}}$ for $m\in\mathbb{Z}$ such that $a_{m}(k)$ takes up the interval $\Omega/2+m\Omega,\Omega/2+m\Omega)$. We summarize this fact in the following lemma, together with Lemma 3.3 and the decay condition (38) in terms of $\alpha$.
@@
where $\alpha$ is defined in (39). The function $\alpha(k)$ can be chosen to be analytic on the intervals $\Omega/2+m\Omega,\Omega/2+m\Omega)$ for $m\in\mathbb{Z}$. Furthermore, $|\absolutevalue{\alpha(k)}$ decays no slower than $|k|^{-2}$ as $|k|\to\infty$.
@@
Since $u(x,k)$ may not satisfy (34) (see Remark 3.2), the function $\alpha$ can be a discontinuous function in $\mathbb{R}$. Lemma 3.3 only ensures $\alpha$ is analytic in each interval $\Omega/2+m\Omega,\Omega/2+m\Omega)$ for $m\in\mathbb{Z}$. Moreover, the function $\alpha$ is formed from a very special choice of coefficients $a_{m}$ since the corresponding $u(x,k)$ is chosen to be analytic in $k$; this does not happen if one computes $u(x,k)$ at each point $k$ independently (see Remark 2.4) so that $u(x,k)$ is an arbitrarily irregular function in $k$.
@@ 3.2 Properties of Wannier functions
[…]
@@
Suppose $u$ in Lemma 3.5 is chosen such that $\alpha$ is integrable. Then the corresponding Wannier function $W_{0}$ in (46) is well defined and is given by the formula
@@
By the Riemann–Lebesgue lemma, $|\absolutevalue{W_{0}(x)}$ goes to zero as $|\absolutevalue{x}\to\infty$. Suppose further that $\alpha^{\prime}$ and $\alpha^{\prime\prime}$ are also integrable. Then the Fourier transform of $W_{0}$, denoted by $\widehat{W}_{0}$, is given by the formula
@@
Thus, if $\alpha$ is chosen to have poor regularity, the Wannier function $W_{0}$ is poorly localized. To define its variance $\alpha$ needs to be smooth so that $|\absolutevalue{W_{0}(x)}$ goes to zero faster than $|1/\absolutevalue{x}$ as $|\absolutevalue{x}\to\infty$. As a result, we also assume that both $\alpha^{\prime}$ and $\alpha^{\prime\prime}$ are integrable. The following lemma contains the formulas for the first and second moments of $W_{0}$ in terms of $\alpha$ for defining its variance. They are consequences of the one-dimensional case of (7) and (8) in [16], first derived in [2].
@@
Suppose that $u$ in Lemma 3.5 is chosen such that the function $\alpha$ has integrable derivatives $\alpha^{\prime}$ and $\alpha^{\prime\prime}$. We have the following formulas for the first and second moments of $W_{0}$: